A man of mass 70kg climbed up a stair case of height 10m in 20seconds. Calculate it's power. (g=10m/s²)


1
Expert's answer
2021-06-13T07:09:58-0400

The power is equal to the force times speed:


"P = Fv"

The force the man acts is equal to his weigth (assuming it moves with the constant speed):


"F = mg = 70kg\cdot 10m/s^2 = 700N"

where "m = 70kg".

By definition, the speed is:


"v=\dfrac{10m}{20s} = 0.5m/s"

Thus, obtain:


"P = 700N\cdot 0.5m/s = 350W"

ANswer. 350W.